OpenStudy (anonymous):

What is the center and radius of the circle with the equation (x + 10)2 + (y + 15)2 = 121

OpenStudy (anonymous):

center is (-10,-15) radius sqrt(121) or 11

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

center is where (x+10)=0, (y+15)=0 radius is sqrt(121)

OpenStudy (anonymous):

ohhh okay thanks!!

OpenStudy (anonymous):

your welcome